The camera melissa wanted for her birthday is on sale at 28% off the usual price. The amount of the discount is $56.00. What was the oringinal price of the camera?
Answer by Alan3354(69443)

28% = $56
$56 = 0.28*OP
OP = 56/0.28 = $200